At the beginning of the year, manufacturing overhead for the year was estimated to be $802,125. At the end of the year, actual direct labor-hours for the year were 36,270 hours, the actual manufacturing overhead for the year was $775,000, and manufacturing overhead for the year was overapplied by $41,075. If the predetermined overhead rate is based on direct labor-hours, then the estimated direct labor-hours at the beginning of the year used in the predetermined overhead rate must have been: (Do not round intermediate calculation.)
Multiple Choice
34,444 direct labor-hours
35,650 direct labor-hours
36,270 direct labor-hours
32,619 direct labor-hours
Answer #1

Answer : 35,650 Direct Labr Hours

Applied Overheads = actual overhead +Overapplied overheads

applied overhead =775,000+41,075 =816,075

Predetermined overhead rate = Applied overhead / Actual Labor hours = 816,075 / 36,270 = $22,.50

Direct labor Hours = Estiamted Overhead / Predetermiend overhead rate

Direct labor Hours =802,125 /22.50 = 35,650 Direct Labor Hours